OpenStudy (anonymous):

dentify the slope and y-intercept of the graph of the equation. y = 3/5x + 1

OpenStudy (texaschic101):

In y = mx + b form...the slope is in the m position and the y intercept is in the b position y = mx + b y = (3/5)x + 1 so what number is in the m position...that is the slope and what number is in the b position....that is your y intercept

OpenStudy (anonymous):

So the slope is 3 and the intercept is 1?

OpenStudy (texaschic101):

you got the y intercept right.....try the slope again...it is a fraction

OpenStudy (anonymous):

3/5

OpenStudy (texaschic101):

correct....slope is 3/5 and y int is 1 but do not forget...if the equation was : y = -3/5x - 1.....then the slope would be -3/5 and the y int would be -1 all I am saying is do not forget the signs

OpenStudy (anonymous):

Thanks

OpenStudy (texaschic101):

not a problem :)